OpenStudy (just_one_last_goodbye):

Which shows the expressions in the order they would appear on a number line from least to greatest?

OpenStudy (just_one_last_goodbye):

A. \[\sqrt{11}, \sqrt{5} + 3, \frac{ 11 }{ 7 }, 2^2, 3^2 - \sqrt{20}\]

OpenStudy (just_one_last_goodbye):

B. \[3^2 - \sqrt{20}, \frac{ 11 }{ 7 }, \sqrt{11}, \sqrt{5} + 3, 2^2\]

OpenStudy (just_one_last_goodbye):

C. \[\frac{ 11 }{ 7 }, \sqrt{11}, 2^2, 3^2 - \sqrt{20}, \sqrt{5} + 3\]

OpenStudy (just_one_last_goodbye):

D. \[\sqrt{5} + 3, 3^2 - \sqrt{20}, 2^2, \frac{ 11 }{ 7 }, \sqrt{11}\]

OpenStudy (anonymous):

option C..

OpenStudy (just_one_last_goodbye):

u sure?

OpenStudy (anonymous):

yes...100%